# Control of kinetic opinion dynamics in popularity-adaptive social networks

### v1

@@ 2.1 The controlled evolution of social media contacts
$\partial_{t}\int_{\mathbb{R}+}\varphi(c)h_{\varepsilon}(v,c)\mathrm{d}c=$ $\int_{\mathbb{R}_{+}}\left(\varphi^{\prime}(c)\beta\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)c+\frac{\nu^{2}c^{2}}{2}\varphi^{\prime\prime}(c)\right)h_{\varepsilon}(v,c)\mathrm{d}c$
$+\varepsilon\int_{\mathbb{R_{+}}}\left(\frac{\varepsilon\beta^{2}}{2}\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)^{2}c^{2}\varphi^{\prime\prime}(c)\right)h_{\varepsilon}(v,c)\mathrm{d}c$
$+O(\varepsilon^{2}).$
@@
$\lim_{\varepsilon\to 0}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)=\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)$
and using the notation $h(v,c)$ for the limit of $h_{\varepsilon}(v,c)$ as $\varepsilon\to 0$, we get the following simplified expression for the evolution of the density of the controlled connections
$\partial_{t}\int_{\mathbb{R}+}\varphi(c)h(v,c)\mathrm{d}c=\int_{\mathbb{R}_{+}}\left(\varphi^{\prime}(c)\beta\left(\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\lambda\,\frac{\beta}{\gamma_{c}}\,R_{c}(c)\,H_{c}(v)\right)c+\frac{\nu^{2}c^{2}}{2}\varphi^{\prime\prime}(c)\right)h(v,c)\mathrm{d}c,$
which is the weak form of the following Fokker-Planck equation for $h$
$\partial_{t}h(v,c)=-\partial_{c}\left(\beta h(v,c)\left(\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\lambda\,\frac{\beta}{\gamma_{c}}\,R_{c}(c)\,H_{c}(v)\right)c-\frac{\nu^{2}}{2}\partial_{c}\left(c^{2}h(v,c)\right)\right).$ (10)
@@ 3 The controlled opinion evolution
Consider two interacting agents with pre-interaction opinions $v,v_{*}$ and contact levels $c,c_{*}$. We denote by $v^{\prime},v_{*}^{\prime}$ the corresponding post-interaction opinions, which evolve through binary interactions, which depends on deterministic and stochastic factors, of the form
$\begin{array}[]{ll}v^{\prime}&=v+\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right)+D(v)\xi,\\ v_{*}^{\prime}&=v_{*}+\alpha\left(P(v_{*},v,c_{*},c)\,(v-v_{*})+u_{*}\right)+D(v_{*})\xi_{*}.\end{array}$ (11)
The random part is given by some unpredictable fluctuation in the opinion which is due, for example, to the agent’s access to information through the news, and is modeled by a random variable $\xi$ with mean equal to zero and given finite standard deviation $\sigma$. This variable is weighted by a self diffusion function $D(v)=1-v^{2}$ that is equal to zero at the boundary of the domain [1,1][-1,1].
The deterministic part of the interaction is modeled by the compromise function $P(v,v_{*},c,c_{*})$, which depends on the opinions of the agents interacting and on their number of connections. The strength of this exchange is expected to be stronger when the two opinions $v$ and $v_{*}$ are close to each other, and when the influence associated with one individual’s connectivity exceeds that of the other. This can be mathematically modeled by expressing the compromise function as the product of two functions
@@
$K(c,c_{*})=\frac{c_{*}^{p}}{c^{p}+c_{*}^{p}}$
@@
The other deterministic factor is given by the control input $u$. Here, we denote by $u^{\star}$ the optimal control obtained as the minimizer of the quadratic cost functional defined by
$u^{\star}=\arg\min_{u}\,\mathcal{V}[u],\qquad\mathcal{V}[u]=\mathbb{E}\!\left[\frac{1}{2}\big(v^{\prime}-\tilde{v}\big)^{2}R_{v}(c)H_{v}(v)+\frac{\gamma_{v}}{2}\alpha\,u^{2}\right].$
The functional $\mathcal{V}$ balances two competing effects. The first term measures the deviation of the post-interaction opinion $v^{\prime}$ from a prescribed target value $\tilde{v}$, that we can suppose is depending on the initial distribution, i.e., $\tilde{v}=\tilde{v}(v_{0})$, and is weighted by the activation functions $R_{v}(c)$ and $H_{v}(v)$, which restrict the action of the control to agents that are both sufficiently active in terms of contacts and lie in a relevant opinion region. The second term penalizes the control magnitude through the parameter $\gamma_{v}>0$, ensuring that excessively strong interventions are discouraged.
Overall, the optimal feedback control $u^{\star}$ tends to steer the opinion toward the target one $\tilde{v}$, but only when the activation conditions encoded by $R_{v}(c)$ and $H_{v}(v)$ are satisfied. The choices for such activation functions will be described in Section 4. We are now interested in an explicit expression for the control $u_{i}^{\star}$. First, we derive $\mathcal{V}[u]$ in $u$, noticing that $\partial_{u}v^{\prime}=\alpha$, and we get
$\partial_{u}\mathcal{V}[u]=\mathbb{E}\!\left[\big(v^{\prime}-\tilde{v}\big)\alpha R_{v}(c)H_{v}(v)+\gamma_{v}\alpha u\right]$
Then, imposing $\partial_{u}\mathcal{V}[u]=0$ and simplifying $\alpha>0$, we substitute the expression of $v^{\prime}$ and we pass to the expected value taking into account the fact that the random variable $\xi$ is centered in zero. In the end, we obtain
$(v+\alpha(P(v,v_{*},c,c_{*})(v_{*}-v)+u)-\tilde{v})R_{v}(c)H_{v}(v)+\gamma_{v}u=0.$
@@
$u^{\star}=-\,\frac{R_{v}(c)H_{v}(v)\,[\,v+\alpha\left(P(v,v_{*},c,c_{*})(v_{*}-v)\right)-\tilde{v}\,]}{\gamma_{v}+\alpha R_{v}(c)H_{v}(v)}.$
@@
Also in the opinion case, just like we did for the contacts in Section 2.1, we can assume that the opinion evolution is due to small changes of the opinions happening with a high frequency. To this aim, we scale the opinion exchange intensity using a small positive parameter $\varepsilon$. Suppose agent $i$ has the pre-interaction couple of opinion and contacts given by $(v,c)$, while the agent $j$ has opinion and contacts given by $(v_{*},c_{*})$, the scaled version of the binary interaction (11) reads as
$\begin{cases}v^{\prime}=v+\varepsilon\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right)+D(v)\xi^{\varepsilon},\\ v_{*}^{\prime}=v_{*}+\varepsilon\alpha\left(P(v_{*},v,c_{*},c)\,(v-v_{*})+u_{*}\right)+D(v_{*})\xi^{\varepsilon}_{*},\end{cases}$
with $\xi^{\varepsilon},\xi^{\varepsilon}_{*}$ random variables with mean equal to zero and variance $\mathbb{E}\left[\left(\xi^{\varepsilon}\right)^{2}\right]=\mathbb{E}\left[\left(\xi^{\varepsilon}_{*}\right)^{2}\right]=\varepsilon\sigma^{2}$. The scaled density of opinion and contacts $f_{\varepsilon}(v,c,t)$ verifies the following Boltzmann-type equation
@@
$\qquad\qquad\left.-\varphi(v,c)-\varphi(v_{*},c_{*})\right)f_{\varepsilon}(v,c,t)f_{\varepsilon}(v_{*},c_{*},t)\mathrm{d}v\mathrm{d}v_{*}\mathrm{d}c\mathrm{d}c_{*}\Bigg],$
with $\varphi(v,c)$ smooth test function. In order to write a simplified expression of (12) we compute the Taylor expansions up to order three of both $\varphi(v^{\prime},c^{\prime})$ and $\varphi(v^{\prime}_{*},c^{\prime}_{*})$ around $v$ and $v_{*}$, respectively. To do so, we compute the increments
$\mathbb{E}\left[v^{\prime}-v\right]$ $=\varepsilon\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right),$
$\mathbb{E}\left[\left(v^{\prime}-v\right)^{2}\right]$ $=\varepsilon^{2}\alpha^{2}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right)^{2}+\varepsilon\sigma^{2}D(v)^{2},$
and
$\mathbb{E}\left[\left(v^{\prime}-v\right)^{3}\right]=$ $\varepsilon^{3}\alpha^{3}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right)^{3}+\mathbb{E}\left[\left(D(v)\xi\right)^{3}\right]$
$+3\varepsilon^{2}\alpha\sigma^{2}D(v)^{2}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u\right).$
Taking into consideration also (7), (8), and (9), and using them in (12) we get
$\partial_{t}\int_{\mathbb{R}_{+}}\int_{-1}^{1}f_{\varepsilon}(v,c,t)\varphi(v,c)\mathrm{d}v\mathrm{d}c=$ $\varepsilon\int_{\mathbb{R}_{+}}\int_{-1}^{1}\left(\alpha\partial_{v}\varphi(v,c)\left(\mathcal{P}[f_{\varepsilon}](v,c,t)+u\right)+\right$. (13)
$+\partial_{c}\varphi(v,c)\beta\left(\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\kappa\right)c$
$+\partial_{vv}^{2}\varphi(v,c)\frac{\sigma^{2}D(v)^{2}}{2}$
$\left.+\partial_{cc}^{2}\varphi(v_{*},c_{*})\frac{\nu^{2}c^{2}}{2}\right)f_{\varepsilon}(v,c,t)\mathrm{d}v\mathrm{d}c+\mathcal{R}(\varepsilon),$
where the reminder $\mathcal{R}(\varepsilon)=O(\varepsilon^{2})$ and the nonlocal operator $\mathcal{P}[f_{\varepsilon}]$ is defined as
$\mathcal{P}\left[f_{\varepsilon}\right](v,c,t)=\int_{\mathbb{R}_{+}}\int_{-1}^{1}P(v,v_{*},c,c_{*})(v_{*}-v)f_{\varepsilon}(v_{*},c,t)\mathrm{d}v_{*}\mathrm{d}c_{*}.$
To model the higher frequency of interaction we can scale the time variable using the same scaling parameter $\varepsilon$, leading to the scaling $t\to\varepsilon t$. Now, we let $\varepsilon\to 0$: for the control variable we have that
$\lim_{\varepsilon\to 0}u=-\,\frac{R_{v}(c)H_{v}(\rho(v))\,[\,v-\tilde{v}\,]}{\gamma_{v}},$
and using the notation $f(v,c,t)$ for the limit of $f_{\varepsilon}(v,c,t)$ when $\varepsilon\to 0$ we have that equation (13) tends to
$\partial_{t}\int_{\mathbb{R}_{+}}\int_{-1}^{1}f(v,c,t)\varphi(v,c)\mathrm{d}v\mathrm{d}c=\int_{\mathbb{R}_{+}}\int_{-1}^{1}\left(\alpha\partial_{v}\varphi(v,c)\left(\mathcal{P}[f](v,c,t)+\right.\right.\\ \left.\frac{R_{v}(c)H_{v}(\rho(v))\,\left(\,v-\tilde{v}\,\right)}{\gamma_{v}}\right)+\partial_{c}\varphi(v,c)\beta\left(\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)+\Phi(v)-\lambda\,\frac{\beta}{\gamma_{c}}\,R_{c}(c)\,H_{c}(v)\right)c\\ \left.+\partial_{vv}^{2}\varphi(v,c)\frac{\sigma^{2}D(v)^{2}}{2}+\partial_{cc}^{2}\varphi(v_{*},c_{*})\frac{\nu^{2}c^{2}}{2}\right)f(v,c,t)\mathrm{d}v\mathrm{d}c,$

### v2

@@ 2.1 The controlled evolution of social media contacts
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}+}$ $\varphi(c){\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}}\mathrm{d}c=\int_{\mathbb{R}_{+}}\left(\varphi^{\prime}(c){\color[rgb]{0,0,0}{\left(-\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)c}}\right.\left.+\frac{\nu^{2}c^{2}}{2}\varphi^{\prime\prime}(c)\right){\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}}\mathrm{d}c$
$+{\color[rgb]{0,0,0}{\frac{\varepsilon}{2}}}\int_{\mathbb{R_{+}}}\left({\color[rgb]{0,0,0}{\left(-\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-\beta\left(\Phi(v)-\kappa_{\ell}\right)\right)^{2}c^{2}}}\varphi^{\prime\prime}(c)\right){\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}}\mathrm{d}c+O(\varepsilon^{1/2}).$
@@
${\color[rgb]{0,0,0}{\lim_{\varepsilon\to 0}\frac{1}{\varepsilon}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)}}=\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)$
and using the notation $\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}$} for the limit of $\color[rgb]{0,0,0}{h_{\ell,\varepsilon}(c,t|v)}$} as $\varepsilon\to 0$, we get the following simplified expression for the evolution of the density of the controlled connections
$\frac{\mathrm{d}}{\mathrm{d}t}\int_{\mathbb{R}+}\varphi(c){\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}}\mathrm{d}c=\int_{\mathbb{R}_{+}}\left[\left(\varphi^{\prime}(c)\left(-\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)-{\color[rgb]{0,0,0}{\beta\left(\Phi(v)-\kappa_{\ell}\right)}}\right)c\right.+\frac{\nu^{2}c^{2}}{2}\varphi^{\prime\prime}(c)\right]{\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}}\mathrm{d}c,$
which is the weak form of the following Fokker-Planck equation for $h_{\ell}(c,t|v)$
$\partial_{t}{\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}}=-\partial_{c}\left(\left(-\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)-{\color[rgb]{0,0,0}{\beta\left(\Phi(v)-\kappa_{\ell}\right)}}\right)c{\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}}-\frac{\nu^{2}}{2}\partial_{c}\left(c^{2}{\color[rgb]{0,0,0}{h_{\ell}(c,t|v)}}\right)\right),$ (11)
coupled with zero-flux bloundary conditions in $c=0$.
Remark 2.2 (Conditional equilibria and quasi-stationary behaviour).
In general, equation (11) does not admit an explicit global steady state due to the drift term coupling the contact dynamics with both the opinion field $\Phi(v)$ and the control $\kappa_{\ell}(v,c)$.
However, for any fixed $v$, (11) becomes a Fokker–Planck equation in $c$. Imposing $\partial_{t}h_{\ell}(c,t|v)=0$ and zero flux boundary conditions leads to
$h_{\ell,\infty}(c|v)=\frac{C_{\ell,\infty}(v)}{c^{2}}\exp\!\left\{-\frac{\mu}{2\nu^{2}}\left(\ln\frac{c}{\overline{c}}\right)^{2}+\frac{2\beta}{\nu^{2}}\int^{c}\frac{\kappa_{\ell}(v,s)-\Phi(v)}{s}\,ds\right\},$
with $C_{\ell,\infty}(v)$ fixed by mass conservation.
The uncontrolled mechanism produces a log-normal confinement around $\overline{c}$ as in [5], while the control acts as an effective potential. A persistent positive drift shifts mass toward large connectivity, a negative one toward a vanishing number of contacts, and only when $\kappa_{\ell}(v,c)$ asymptotically balances $\Phi(v)$ then a normalizable stationary state exists.
This means that $h_{\ell,\infty}(c|v)$ represents a conditional (quasi-stationary) equilibrium: the network rapidly relaxes in $c$ given $v$, while the slow opinion dynamics continuously moves the equilibrium. A steady state arises when control and opinion feedback become independent of time and $v$.
@@ 3 The controlled opinion evolution
Let us consider two interacting agents with pre-interaction opinions $v,v_{*}$ and contact levels $c,c_{*}$, respectively, belonging to populations $\ell$ and $k$. Notably, the two agents may belong either to the same population or to two different ones. We denote by $v^{\prime},v_{*}^{\prime}$ the corresponding post-interaction opinions, which evolve through binary interactions depending on deterministic and stochastic factors of the form
$\begin{array}[]{ll}{\color[rgb]{0,0,0}{v^{\prime}}}&{\color[rgb]{0,0,0}{=v+\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+u_{\ell}\right)+D(v)\xi,}}\\ {\color[rgb]{0,0,0}{v_{*}^{\prime}}}&{\color[rgb]{0,0,0}{=v_{*}+\alpha\left(P(v_{*},v,c_{*},c)\,(v-v_{*})+u_{k}\right)+D(v_{*})\xi_{*}.}}\end{array}$ (12)
The random part is given by some unpredictable fluctuation in the opinion which is due, for example, to the agent’s access to information through the news, and is modeled by random variables $\xi$ and $\xi_{*}$, with mean equal to zero, given finite standard deviation $\sigma$, and $\mathbb{E}\left[\lvert\xi\rvert^{3}\right]=\mathbb{E}\left[\lvert\xi_{*}\rvert^{3}\right]<+\infty$. These variables are weighted by a self diffusion function $D(v)=1-v^{2}$ that is equal to zero at the boundary of the domain [1,1][-1,1].
The deterministic part of the interaction is modeled by the compromise function $P(v,v_{*},c,c_{*})$, depending on the opinions of the agents interacting and on their number of connections. The strength of this exchange is expected to be stronger when the two opinions $v$ and $v_{*}$ are close to each other, and when the influence associated with one individual’s connectivity exceeds that of the other. This can be mathematically modeled by expressing the compromise function as the product of two functions
@@
$\color[rgb]{0,0,0}{K(c,c_{*})=\begin{cases}\frac{c_{*}^{p}}{c^{p}+c_{*}^{p}},\quad&\mbox{if }c^{p}+c^{p}_{*}\neq 0,\\ 0,&\mbox{otherwise},\end{cases}}$}
@@
The other deterministic contribution is given by the control input $u_{\ell}$, which models an external intervention acting on the opinion of agents belonging to population $\ell$. As for the contact control, we adopt a common functional form of the opinion control for all populations. The differences among populations are encoded through different choices of the control parameters and of the threshold functions appearing in its definition. We denote by $u_{\ell}^{\star}$ the optimal control obtained as the minimizer of the quadratic cost functional defined by
$\color[rgb]{0,0,0}{u^{\star}_{\ell}=\arg\min_{u}\,\mathcal{V}_{\ell}[u],\qquad\mathcal{V}_{\ell}[u]=\mathbb{E}\!\left[\frac{1}{2}\big(v^{\prime}-\tilde{v}_{\ell}\big)^{2}R^{\ell}_{v}(c)H^{\ell}_{v}(v)+\frac{\gamma_{v}^{\ell}}{2}\alpha\,u^{2}\right]}$}.
The functional $\mathcal{V}_{\ell}$ balances two competing effects. The first term measures the deviation of the post-interaction opinion $v^{\prime}$ from a prescribed target value $\tilde{v}_{\ell}$, different for each population $\ell$, and is weighted by the activation functions $R^{\ell}_{v}(c)$ and $H^{\ell}_{v}(v)$, which restrict the action of the control to agents that are both sufficiently active in terms of contacts and lie in a relevant opinion region. The second term penalizes the control magnitude through the parameter $\gamma^{\ell}_{v}>0$, ensuring that excessively strong interventions are discouraged.
Overall, the optimal feedback control $u^{\star}_{\ell}$ aims to steer the opinion toward the target one $\tilde{v}_{\ell}$, but only when the activation conditions encoded by $R^{\ell}_{v}(c)$ and $H^{\ell}_{v}(v)$ are satisfied. The specific choices for such activation functions will be detailed in the numerical tests described in Section 4. We are now interested in an explicit expression for the control $\color[rgb]{0,0,0}{u_{\ell}^{\star}}$}. First, we derive $\mathcal{V}_{\ell}[u]$ in $u$, noticing that $\partial_{u}v^{\prime}=\alpha$, and we get
$\color[rgb]{0,0,0}{\partial_{u}\mathcal{V}_{\ell}[u]=\mathbb{E}\!\left[\big(v^{\prime}-\tilde{v}_{\ell}\big)\alpha R^{\ell}_{v}(c)H^{\ell}_{v}(v)+\gamma^{\ell}_{v}\alpha u\right]}$}.
Then, imposing $\partial_{u}\mathcal{V}_{\ell}[u]=0$ and simplifying $\alpha>0$, we substitute the expression of $v^{\prime}$ and we pass to the expected value taking into account the fact that the random variable $\xi$ is centered in zero. In the end, we obtain
$\color[rgb]{0,0,0}{(v+\alpha(P(v,v_{*},c,c_{*})(v_{*}-v)+u)-\tilde{v}_{\ell})R^{\ell}_{v}(c)H^{\ell}_{v}(v)+\gamma^{\ell}_{v}u=0.}$}
@@
$\color[rgb]{0,0,0}{u_{\ell}^{\star}=-\,\frac{R^{\ell}_{v}(c)H^{\ell}_{v}(v)\,\left(\,v+\alpha\left(P(v,v_{*},c,c_{*})(v_{*}-v)\right)-\tilde{v}_{\ell}\,\right)}{\gamma_{v}^{\ell}+\alpha R^{\ell}_{v}(c)H^{\ell}_{v}(v)}.}$} (13)
@@
The evolution of the density of the $\ell$-th population of agents $f_{\ell}(v,c,t)$ is described by the Boltzmann-like equation
$\color[rgb]{0,0,0}{\partial_{t}f_{\ell}(v,c,t)=\sum_{k}Q_{\ell}(f_{\ell},f_{k})(v,c,t)}$},
where the Boltzmann operators $Q_{\ell}$ between populations $\ell$ and $k$, in weak form, read as
${\color[rgb]{0,0,0}{\iint_{[-1,1]\times\mathbb{R}_{+}}\varphi(v,c)Q_{\ell}(f_{\ell},f_{k})\mathrm{d}v\mathrm{d}c=}}\\ {\color[rgb]{0,0,0}{\iiiint_{[-1,1]^{2}\times\mathbb{R}^{2}}\mathbb{E}\left[\varphi(v^{\prime},c^{\prime})-\varphi(v,c)\right]f_{\ell}(v,c,t)f_{k}(v_{*},c_{*},t)\mathrm{d}v\mathrm{d}v_{*}\mathrm{d}c\mathrm{d}c_{*},}}$ (14)
with $\varphi(v,c)$ smooth test function. Also in the opinion case, just like we did for the contacts in Section 2.1, we can assume that the opinion evolution is due to small changes of the opinions happening with a high frequency. To this aim, we scale the opinion exchange intensity using the quasi-invariant scaling $\alpha\to\varepsilon\alpha$ and $\sigma^{2}\to\varepsilon\sigma^{2}$ for a small positive parameter $\varepsilon$, therefore the scaled version of the binary interaction (12) reads as
$\begin{cases}v^{\prime}=v+\varepsilon\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+{\color[rgb]{0,0,0}{u_{\ell,\varepsilon}}}\right)+D(v)\xi^{\varepsilon},\\ v_{*}^{\prime}=v_{*}+\varepsilon\alpha\left(P(v_{*},v,c_{*},c)\,(v-v_{*})+{\color[rgb]{0,0,0}{u_{k,\varepsilon}}}\right)+D(v_{*})\xi^{\varepsilon}_{*},\end{cases}$
with $\xi^{\varepsilon},\xi^{\varepsilon}_{*}$ random variables with mean equal to zero, variance $\mathbb{E}\left[\left(\xi^{\varepsilon}\right)^{2}\right]=\mathbb{E}\left[\left(\xi^{\varepsilon}_{*}\right)^{2}\right]=\varepsilon\sigma^{2}$, and $\mathbb{E}\left[\left(\xi^{\varepsilon}\right)^{3}\right]=\mathbb{E}\left[\left(\xi_{*}^{\varepsilon}\right)^{3}\right]=O(\varepsilon^{3/2})$. For each population $\ell$, the scaled density of opinion and contacts $f_{\ell,\varepsilon}(v,c,t)$ verifies the following Boltzmann-type equation
$\color[rgb]{0,0,0}{\partial_{t}f_{\ell,\varepsilon}(v,c,t)=\sum_{k}Q_{\ell}(f_{\ell,\varepsilon},f_{k,\varepsilon})(v,c,t)}$}, (15)
where the operators $Q_{\ell}$ in (15) are defined as in (14).
In order to write a simplified expression of (15) we compute the Taylor expansions up to order three of both $\varphi(v^{\prime},c^{\prime})$ and $\varphi(v^{\prime}_{*},c^{\prime}_{*})$ around $v$ and $v_{*}$, respectively. To do so, we compute the increments
$\mathbb{E}\left[v^{\prime}-v\right]$ $=\varepsilon\alpha\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+{\color[rgb]{0,0,0}{u_{\ell,\varepsilon}}}\right),$
$\mathbb{E}\left[\left(v^{\prime}-v\right)^{2}\right]$ $=\varepsilon^{2}\alpha^{2}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+{\color[rgb]{0,0,0}{u_{\ell,\varepsilon}}}\right)^{2}+\varepsilon\sigma^{2}D(v)^{2},$
and
$\mathbb{E}\left[\left(v^{\prime}-v\right)^{3}\right]=$ $\varepsilon^{3}\alpha^{3}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+{\color[rgb]{0,0,0}{u_{\ell,\varepsilon}}}\right)^{3}+\mathbb{E}\left[\left(D(v)\xi{\color[rgb]{0,0,0}{{}^{\varepsilon}}}\right)^{3}\right]$
$+3\varepsilon^{2}\alpha\sigma^{2}D(v)^{2}\left(P(v,v_{*},c,c_{*})\,(v_{*}-v)+{\color[rgb]{0,0,0}{u_{\ell,\varepsilon}}}\right).$
Taking into consideration also (8), (9), and (10), and using them in (15) , we obtain, after summing over $k$,
$\frac{\mathrm{d}}{\mathrm{d}t}\iint_{[-1,1]\times\mathbb{R}_{+}}{\color[rgb]{0,0,0}{f_{\ell,\varepsilon}}}(v,c,t)\varphi(v,c)\mathrm{d}v\mathrm{d}c=$ (16)
$\,\varepsilon\iint_{[-1,1]\times\mathbb{R}_{+}}\left(\alpha\partial_{v}\varphi(v,c)\left({\color[rgb]{0,0,0}{\sum_{k}\mathcal{U}_{\ell,\varepsilon}[f_{k,\varepsilon}](v,c,t)}}\right)\right$.
$\,\,+\partial_{c}\varphi(v,c)\left({\color[rgb]{0,0,0}{-\frac{1}{\varepsilon}}}\Psi^{\varepsilon}\left(\frac{c}{\overline{c}}\right)-{\color[rgb]{0,0,0}{\beta\left(\Phi(v)-\kappa_{\ell}\right)}}\right)c$
$\,\,\,\left.+\partial_{vv}^{2}\varphi(v,c)\frac{\sigma^{2}D(v)^{2}}{2}+\partial_{cc}^{2}\varphi(v,c)\frac{\nu^{2}c^{2}}{2}\right){\color[rgb]{0,0,0}{f_{\ell,\varepsilon}}}(v,c,t)\mathrm{d}v\mathrm{d}c+\mathcal{R}(\varepsilon),$
@@
$\color[rgb]{0,0,0}{\mathcal{U}_{\ell,\varepsilon}\left[f_{k,\varepsilon}\right](v,c,t)=\iint_{[-1,1]\times\mathbb{R}_{+}}\left(P(v,v_{*},c,c_{*})(v_{*}-v)+u_{\ell,\varepsilon}\right)f_{k,\varepsilon}(v_{*},c_{*},t)\mathrm{d}v_{*}\mathrm{d}c_{*}}$}.
We stress that the contact control $\kappa_{\ell}=\kappa_{\ell}(v,c)$ depends only on the local state variables $(v,c)$ of the tagged agent, whereas the opinion control $u_{\ell}=u_{\ell}(v,c,v_{*},c_{*})$ (and its rescaled version) may in general depend on the states of both interacting agents. At the kinetic level, this distinction is reflected in the limit equation: the former enters as a local drift term acting on $f_{\ell}$, while the latter contributes to the nonlocal drift operator $\mathcal{U}_{\ell,\varepsilon}[f_{k}]$ through its dependence on the interacting density.
To model the higher frequency of interaction we can scale the time variable using the same scaling parameter $\varepsilon$, leading to the scaling $t\to\varepsilon t$.
Now, we let $\varepsilon\to 0$. We assume that the rescaled control variable admits a finite limit denoted as $\bar{u}_{\ell}:=\lim_{\varepsilon\to 0}u_{\ell,\varepsilon}$, and using the notation $\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}$} for the limit of $\color[rgb]{0,0,0}{f_{\ell,\varepsilon}(v,c,t)}$} we have that equation (16) tends to
$\frac{\mathrm{d}}{\mathrm{d}t}\iint_{[-1,1]\times\mathbb{R}_{+}}{\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\varphi(v,c)\mathrm{d}v\mathrm{d}c=\int_{\mathbb{R}_{+}}\int_{-1}^{1}\left(\alpha\partial_{v}\varphi(v,c)\left({\color[rgb]{0,0,0}{\sum_{k}\mathcal{U}_{\ell}[f_{k}](v,c,t)}}\right)\right.\\ +\partial_{c}\varphi(v,c)\left(-\frac{\mu}{2}\ln\left(\frac{c}{\overline{c}}\right)-{\color[rgb]{0,0,0}{\beta\left(\Phi(v)-\kappa_{\ell}\right)}}\right)c\\ \left.+\partial_{vv}^{2}\varphi(v,c)\frac{\sigma^{2}D(v)^{2}}{2}+\partial_{cc}^{2}\varphi(v,c)\frac{\nu^{2}c^{2}}{2}\right){\color[rgb]{0,0,0}{f_{\ell}(v,c,t)}}\mathrm{d}v\mathrm{d}c,$
with
$\color[rgb]{0,0,0}{\mathcal{U}_{\ell}\left[f_{k}\right](v,c,t)=\iint_{[-1,1]\times\mathbb{R}_{+}}\left(P(v,v_{*},c,c_{*})(v_{*}-v)+\bar{u}_{\ell}\right)f_{k}(v_{*},c_{*},t)\mathrm{d}v_{*}\mathrm{d}c_{*}}$}, (17)